OpenStudy (anonymous):

Solve C = 3(2f – 5h) for f in terms of C and h. You must show all work

OpenStudy (anonymous):

C=6f-15h or C-15h = 6f or f= [C-15h]/6

OpenStudy (anonymous):

what i did was i distributed the 3 first got 6f - 15h = C then i added 15h to both sides 6f = C+15h then divided both sides by 6 thats correct right ?

OpenStudy (anonymous):

yes yes ... you are right C=6f-15h or C+15h = 6f or f= [C+15h]/6 should be the answer ... had made a careless mistake earlier..sorry about that !!